OpenStudy (anonymous):

Divide. 4/5 divided 2/3 A.8/15 B.5/6 C.1and1/5 D.1and7/8

OpenStudy (agl202):

6/5 = 1 1/5

OpenStudy (anonymous):

\[\huge \frac{ a }{ b }\div \frac{ c }{ d }=\frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

Truly the value of learning is lost when the answer is given without the methodology to solving. I suggest explaining how you got your answer next time @Agl202